User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Lucas

next week's chair: Lim

Attended: Ross, Jamie, Simin, Eloise, Earl, Mas, Jay, Becca, Katie , Patrick

Apologies: Prad

Absent: Lim

General Issues

  • Shifts
    • Back to back overnight shifts?
    • Day shifts
    • Warren away currently for 2 weeks
  • 12m VGOS experiments starting this week - mixed mode.
    • Fringes to Ke-Ho, Ke-Hb. Correlated Vienna.
    • Setup/monitoring notes are partially available, will be updated soon.
    • 26th Feb.
  • Hobart
    • One Hobart26m drive controller replaced, ongoing process but better now. Needs pointing update (lots of slewing errors). New pointing solution required.
    • jive5ab might be more robust? Throws errors though (benign)
  • Training shifts?
    • Next observers meeting for mixed mode training for all observers.

Handover Issues

  • R1930
    • Hobart26m multiple drive crashes.
  • AUA060
    • Hobart26m Faulty m5khb disk?
      • Power cycling, resyncing didn't initially work, fmset etc. Disk pack most likely the problem. If so disk needs replacing.
/home/www/auscope/opswiki/data/pages/operations/agenda2019_01_30.txt · Last modified: 2020/01/30 02:15 by Tiege McCarthy